OpenStudy (anonymous):

A manufacturer bought a new rolling press for 48,000. It has depreciated in value at an annual rate of 15%. What is its value 5 yrs after purchase? Round to the nearest hundred dollars.

OpenStudy (tkhunny):

What is left after 15% is gone?

OpenStudy (anonymous):

im not sure please help

OpenStudy (tkhunny):

Let's try this. You had 100 sheep, but 15% of them ran off and were eaten by coyotes. How many sheep remain in your care?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

85?

OpenStudy (tkhunny):

There you go. You DO have an idea. Never say again that you don't. 15% of 100 is 15 100 * 15% = 100 * 0.15 = 15 After 15% goes away, we have: 100% - 15% = 85% Start: 48000 1 Year Depreciation: 48000*0.85 = 40,800 2 Years Depreciation: 40800*0.85 = 34,680 See how that works?